KPSEWHICH(1) KPSEWHICH(1) NAME kpsewhich−standalonepathlookupandandexpansionforkpathsea SYNOPSIS kpsewhich[options][filenames] DESCRIPTION Thismanualpageisnotmeanttobeexhaustive.Thecompletedocumentationforthisversionof TEXcanbefoundintheinfofileormanualKpathsea:Alibraryforpathsearching. kpsewhichisusedasastandalonefront-endofthekpathsealibrarythatcanbeusedtoexamine variablesandfindfiles.Whenthe--formatoptionisnotgiven,thesearchpathusedwhenlook-ingforafileisinferredfromthenamegiven,bylookingforaknownextension.Ifnoknown extensionisfound,thesearchpathforTEXsourcefilesisused. OPTIONS kpsewhichacceptsthefollowingoptions: --debugnum Setdebuggingflags. -Dnum Useabaseresolutionofnum;thedefault,setbytheinstaller,istypically600. --dpinum As-D. --expand-bracesstring Printvariableandbraceexpansionofstring. --expand-pathstring Printcompletepathexpansionofstring. --expand-varstring Printvariableexpansionofstring. --formatname Usefiletypename.Seetheinfomanualforalistofvalidnames,orusethe--help optiontoprintthelist. --helpPrinthelpmessageandexit. --interactive Askforadditionalfilenamestolookup. --mktexfmt enablemktexfmtgeneration.(fmt=pk/mf/tex/tfm) --modestring Setdevicenamefor$MAKETEXMODEtostring;nodefault. --must-exist Searchthediskaswellasls-Rifnecessary. --no-mktexfmt disablemktexfmtgeneration.(fmt=pk/mf/tex/tfm) --pathstring Searchinthepathstring. --prognamestring Setprogramnametostring. --show-pathname Outputsearchpathforfiletypename.Seetheinfomanualforalistofvalidnames,or usethe--helpoptiontoprintthelist. Kpathsea7.32x 4January1998 1 KPSEWHICH(1) KPSEWHICH(1) --version Printversioninformationandexit. SEEALSO mktexlsr(1),mktexmf(1),mktexpk(1),mktextfm(1). Kpathsea7.32x 4January1998 2